OpenStudy (anonymous):

a circular cylinder with a volume of 8π m^3 is circumscribed about a right prism whose base is an equilateral triangle with one side that measures 2 m . What is the altitude of the cylinder?

OpenStudy (anonymous):

8 pi *

OpenStudy (anonymous):

@CGGURUMANJUNATH

OpenStudy (anonymous):

 tan 30=(sq.rt3)/3= x/1 r^2=x^2+1^2 3/9 +1=12/9 =4/3 vol cylinder= pi*r^2*h=pi*8 given r^2h=8 4/3*h=8 h=8 *3/4 h=6 hope this helps